OpenStudy (anonymous):

Solve the system using elimination. 2x+6y=-12 5x-5y =10

OpenStudy (anonymous):

7x+1y=-2 ?!?!?!?!?!?!?!? Feels right to me.

OpenStudy (tkhunny):

?? Nothing was eliminated. Multiply the first by 5. Multiply the second by -2. Add the two. x is eliminated.

OpenStudy (anonymous):

??????

OpenStudy (mew55):

i'll try to explain okay, letz get rid of the x value. we will get rid of it by multiplying the first equation by 5 so u can get 5(2x+6y=-12)=10x+30y=-60 Second, letz multiply the second equation by -2 -2(5x-5y=10)=-10x+10y=-20 c how -10+10=0?. The x is cancelled. so u r left w/ 40y=-80 y=-20. But,we r not done yet. we need to find the x value as well so with the answer we got, plug it in the y into any equation. letz do the second equation 5x-5(40)=10 5x=200 x=45 .....did i do it right XD?

OpenStudy (tkhunny):

2x+6y=-12 5x-5y =10 Multiply the first by 5. Multiply the second by -2. 10x+30y=-60 -10x+10y =-20 Add the two eequations.

OpenStudy (mew55):

well is it right XD?

OpenStudy (tkhunny):

Not quite. 40y=-80 y = -2

OpenStudy (mew55):

omg oops XD. im sorry i totally forgot to divide XD

OpenStudy (mew55):

sorry not divide, divided wrong, im sorry XD. okay letz try the step again Revised XD letz plug in -2 in the y. with the second equation 5x-5y=10 5x-5(-2)=10 5x+10=10 5x=0 x=0

OpenStudy (tkhunny):

That's better.

OpenStudy (mew55):

hahahaha XD